Every year the Earth travels about 109 km as it orbitsthe Sun. What is Earth's average speed in km/h?
Answer #1
1 year=365 days
1 day=24 hrs
1 year=365*24 hr

distance traveled/time=average speed
average speed=109/(365*24)=114155.25 km/h